# FusionBench: A Comprehensive Benchmark of Deep Model Fusion

### v3

@@ Appendix A Evaluation of Fine-Tuned Single-Task Models
Figure 5: Cosine similarity matrices of task vectors for CLIP-ViT-B/32 and CLIP-ViT-L/14 models.
Table 9: Performance of fine-tuned single-task GPT-2 models on the seven text classification tasks.
@@
Based on the performance metrics detailed in these tables, we observe that the fine-tuned models demonstrate high accuracy on specific tasks. This observation holds true across various model architectures and task domains, indicating the effectiveness of the fine-tuning process in adapting pre-trained models to excel in particular applications.
What’s more, fine-tuning a model on one task can lead to both positive and negative transfer effects on other tasks. Positive transfer occurs when the knowledge gained from fine-tuning on one task enhances the model’s performance on another task, while negative transfer arises when the fine-tuning process on one task hinders the model’s ability to perform well on other tasks.
Appendix B Multi-Task Model Fusion
@@
Table 12: Multi-task performance when merging GPT-2 models on seven text classification tasks.
@@
Multi-Task Model Fusion Method
Simple Average [57] 55.0 55.1 51.0 57.6 76.7 44.8 52.5 56.1
Fisher Merging [37] 54.8 58.0 39.5 63.3 81.5 49.1 64.7 58.7
RegMean [28] 61.7 70.4 65.4 69.7 78.8 56.0 79.7 68.8
Task Arithmetic [25] 68.7 68.6 69.6 70.5 81.8 47.3 83.6 70.0
Ties-Merging [62] 68.4 71.4 68.4 69.6 82.4 47.7 81.8 70.0
Table 13: Experimental results of merging Flan-T5-base (LoRA fine-tuned) models on all eight tasks.
@@
Weight Averaging [57] 69.7 59.7 78.9 90.1 83.8 80.5 91.2 72.0 78.2
Task Arithmetic [25] 68.8 55.2 78.7 89.8 83.7 79.1 91.5 72.4 77.4
Ties-Merging [62] 68.3 56.3 79.4 89.8 83.7 79.4 91.6 71.2 77.5
Layer-wise AdaMerging [63] 69.1 60.3 78.4 90.0 83.6 79.1 91.6 74.1 78.3
Table 14: Multi-task performance when merging CLIP-ViT-L/14 models on all eight tasks.
Method SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Avg.
Reference Methods
Pre-trained 68.3 77.8 71.0 58.9 58.4 50.6 76.4 55.5 64.6
Fine-tuned (STL) 82.8 92.9 97.4 99.2 97.9 99.2 99.8 85.5 94.3
Multi-Task Model Fusion Methods
Weight Averaging [57] 72.5 81.5 82.2 90.0 81.6 74.0 96.6 61.8 80.0
Fisher Merging [37] 70.6 79.4 84.1 98.1 74.7 85.0 89.5 61.0 80.3
RegMean [28] 75.3 88.4 90.0 97.1 95.9 92.4 98.5 72.6 88.8
Task Arithmetic [25] 72.0 79.0 80.5 86.0 87.5 83.5 98.0 58.8 80.7
Ties-Merging [62] 74.7 83.3 86.4 91.3 89.7 85.2 97.8 63.9 84.0
task-wise AdaMerging [63] 75.8 80.1 77.2 83.6 68.4 93.5 93.1 69.0 80.1
layer-wise AdaMerging [63] 78.1 90.7 90.8 96.5 94.8 97.5 98.6 81.3 91.0
WEMoE (Model Mixing) [51] 81.5 92.3 96.5 98.8 97.6 99.4 99.6 84.5 93.8
@@
Weight Averaging [57] 74.6 84.3 84.1 92.8 86.3 87.4 94.8 88.0 86.5
Task Arithmetic [25] 76.9 85.4 85.3 93.9 85.8 88.1 95.2 87.8 87.3
Ties-Merging [62] 77.1 85.1 86.3 93.9 86.0 87.7 95.1 88.0 87.4
Layer-wise AdaMerging [63] 76.7 87.6 84.8 93.8 85.9 88.1 95.2 88.6 87.6
We begin by comparing the performance of multi-task model fusion algorithms on various tasks using different models. These experiments provide insights into the effectiveness of different fusion methods in improving the performance of multi-task models. We evaluate the performance of multi-task model fusion algorithms on image classification tasks using CLIP models, scene understanding tasks using ResNet-50 models, text classification tasks using GPT-2 models, and text-to-text generation tasks using Flan-T5 models.
  1. 1.
    Image Classification Tasks with CLIP Models: We utilize the CLIP-ViT-B/32 and CLIP-ViT-L/14 models from the HuggingFace Library [24]. The results of merging CLIP-ViT-B/32 models on all eight tasks are provided in Table 3 and Figure 4. The results of CLIP-ViT-L/14 models are shown in Table 14.
  2. 2.
    Scene Understanding Tasks with ResNet-50 Models: We use the NYUv2 dataset and ResNet-50 models for segmentation, depth estimation, and normal estimation tasks. In Figure 6, we illustrate the process of merging ResNet-50 models on these tasks, where the backbones are merged, and the heads are copied separately. The results of merging ResNet-50 models on these tasks are shown in Table 4.
  3. 3.
    Text Classification Tasks with GPT-2 Models: The results of merging GPT-2 models on seven text classification tasks are shown in Table 12.
  4. 4.
    Text-to-Text Generation Tasks with Flan-T5 Models: For LoRA fine-tuned Flan-T5-base and Flan-T5-large models, we merge and unload the LoRA adapters before performing multi-task model fusion. The results of merging Flan-T5-base and Flan-T5-large models on all eight tasks are shown in Tables 13 and 15, respectively.
In the above mentioned tables, we compare the performance of different multi-task model fusion algorithms on various tasks. The results of pre-trained models, fine-tuned models, and traditional multi-task learning (MTL) are provided as reference methods.
Appendix C Generalization Experiments
@@
Table 16: Generalization results on two unseen tasks when merging ViT-B/32 models on six tasks.
Method Seen Tasks Unseen Tasks
@@
Fisher Merging [37] 68.1 67.4 67.2 86.4 58.6 81.6 71.5 60.2 42.5 51.3
RegMean [28] 69.4 70.5 86.9 97.0 67.1 98.3 81.5 50.2 51.5 50.8
Task Arithmetic [25] 65.2 63.6 76.1 87.1 56.4 94.2 73.8 52.4 45.2 48.8
Ties-Merging [62] 68.2 65.9 70.0 81.2 56.0 89.0 71.7 60.3 47.3 53.8
AdaMerging [63] 69.8 72.4 95.5 95.1 70.7 98.1 83.6 48.7 60.7 54.7
WEMoE [51] 74.3 78.1 98.8 98.7 75.1 99.5 87.4 47.3 51.3 49.3
For the generalization experiments, we assess the performance of multi-task model fusion algorithms on two unseen tasks after merging ViT-B/32 models trained on six tasks. The performance of various multi-task model fusion methods, including Fisher Merging [37], RegMean [28], Task Arithmetic [25], Ties-Merging [62], AdaMerging [63], and WEMoE [51], is compared across both the seen tasks and unseen tasks.
Specifically, we conduct two sets of generalization experiments using the CLIP-ViT-B/32 models:

### v4

@@ D.1 CLIP-ViT Models
Figure 4: Cosine similarity of task vectors for CLIP-ViT-B/32 and CLIP-ViT-L/14 models.
The performance of fine-tuned CLIP-ViT-B/32 and CLIP-ViT-L/14 models on the eight most commonly used image classification tasks is shown in Tables 2 and 3, respectively. In Figure 4, we visualize the cosine similarity matrices of task vectors for CLIP-ViT-B/32 and CLIP-ViT-L/14 models. Note that while we present results for the eight core tasks in this paper for conciseness, comprehensive evaluation results for all 20 tasks across CLIP-ViT-B/32, CLIP-ViT-B/16, and CLIP-ViT-L/14 models are available in our online documentation. Following (ilharco2022editing), the task vector is defined as the difference between a fine-tuned model’s parameters and the original pre-trained model’s parameters: $\tau=\theta_{ft}-\theta_{pre}$, where $\theta_{ft}$ represents the fine-tuned weights and $\theta_{pre}$ represents the pre-trained weights. We note that the task vectors for models from various tasks are nearly orthogonal. This suggests that the knowledge specific to each task resides in distinct directions or subspaces. This finding motivates the exploration of locating subspaces in which the knowledge of different tasks can be merged effectively, as discussed in tang2023concrete; wang2024localizing.
D.2 ResNet-50 Models
Table 4: Single-task performance of fine-tuned ResNet-50 models on NYUv2 dataset.
Method SEGMENTATION DEPTH ESTIMATION NORMAL
Segmentation 52.0 73.8 242.8 88.7 82.8
Depth Estimation 2.3 6.2 42.5 17.7 82.8
Normal 2.0 4.9 264.0 98.1 24.7
We fine-tune ResNet-50 models on three distinct scene understanding tasks using the NYUv2 dataset: semantic segmentation, depth estimation, and surface normal estimation. Each model is trained with a learning rate of 1e-4 for 40 epochs, with the learning rate reduced by a factor of 0.5 every 10 epochs. Table 4 presents the single-task performance of these specialized models, where each row represents a model trained exclusively on one task and evaluated across all three tasks using task-specific metrics. The results demonstrate clear task specialization: the segmentation model achieves optimal performance on segmentation metrics (52.0 mIoU, 73.8 Pixel Accuracy), the depth estimation model excels at depth-related metrics (42.5 Absolute Error, 17.7 Relative Error), and the normal estimation model performs best on surface normal prediction (24.7 Mean angular error). Notably, models trained on one task show severely poor performance on other tasks, highlighting the challenge of multi-task learning in computer vision applications.
D.3 GPT-2 Models
Table 5: Performance of fine-tuned GPT-2 models on the seven text classification tasks.
@@ D.3 GPT-2 Models
We evaluate GPT-2 models fine-tuned on seven distinct text classification tasks from the GLUE benchmark, as detailed in the enumeration above. Table 5 presents the cross-task performance evaluation, where each row represents a model fine-tuned on a specific task (indicated by the row label) and columns show its performance across all seven tasks. The results demonstrate clear task specialization, with each fine-tuned model achieving its highest performance on its target task (diagonal elements).
D.4 Flan-T5 Models
@@
We evaluate Flan-T5-base and Flan-T5-large models fine-tuned using LoRA on eight text-to-text generation tasks from the GLUE benchmark (wang2018glue). Tables 6 and 7 present the cross-task performance evaluation for both model sizes, where each row represents a model fine-tuned on a specific task and columns show performance across all eight tasks. The results demonstrate that each specialized model achieves its best performance on its target task (diagonal elements), with the Flan-T5-large models generally outperforming their base counterparts across all tasks. Notably, when evaluating fine-tuned models on non-target tasks, we observe both positive and negative transfer effects. For instance, in Flan-T5-base models, the MNLI-specialized model shows positive transfer to QNLI (89.3% vs 88.% pre-trained baseline), while the STSB-specialized model exhibits negative transfer to MNLI (1.7% vs 56.5% pre-trained baseline).
Appendix E Experimental Evaluation and Benchmarking Results
In this section, we evaluate the performance of multi-task deep model fusion algorithms on a variety of tasks, as well as analyze the generalization and robustness of these algorithms. We also provide an ablation study to investigate the impact of hyperparameter selection. The majority of experiments presented in this section can be executed using a single NVIDIA RTX 3090 GPU with 24GB memory, ensuring cost-effective reproducibility.
E.1 Experimental Setup
In this section, we conduct a series of multi-task deep model fusion experiments on image classification tasks, text classification tasks, and text-to-text generation tasks to evaluate the performance of multi-task deep model fusion algorithms. These tasks are chosen to cover a wide range of NLP and CV tasks, as described in Section E.3.
E.2 Computational Cost Analysis
While functional requirements listed in Table 1 differ across methods, computational resource implications also vary significantly. For instance, ensemble methods incur high inference costs since multiple models must be stored and executed simultaneously. Model merging methods are generally cost-efficient during inference, as the merged model retains the footprint of a single model, though some approaches require additional data-dependent computations during merging. Mixing methods typically expand the model size and necessitate partial re-training, leading to substantial computational overhead.
Practical Implications: For resource-constrained environments, parameter-free methods like parameter average and task arithmetic offer the best cost-performance trade-off. When additional data is available, alignment-based methods such as RegMean (jin2022dataless) and RegMean++ (nguyen2025regmean++) provide strong performance with moderate computational overhead. For applications requiring maximum performance and having sufficient computational resources or access to the test data distribution, test-time adaptive methods like AdaMerging (yang2023adamerging) or model mixing approaches like WE-MoE (tang2024merging) are recommended despite their higher computational costs. Besides, the LazyStateDict utility in FusionBench significantly reduces memory requirements for all methods by enabling on-demand parameter loading, making large-scale model fusion practical.
E.3 Multi-Task Deep Model Fusion
We start our evaluation by comparing multi-task model fusion algorithms across multiple experimental settings, including both computer vision and natural language processing tasks:
  1. 1.
    Open-Vocabulary Image Classification: Using the CLIP-ViT models fine-tuned on the eight most commonly used tasks. Results for CLIP-ViT-B/32 and CLIP-ViT-L/14 models are presented in Section E.3.1.
  2. 2.
    Scene Understanding: Using ResNet-50 models on the NYUv2 dataset for segmentation, depth estimation, and normal estimation tasks. Detailed results are provided in Section E.3.2.
  3. 3.
    Text Classification: Detailed results for GPT-2 models on seven text classification tasks are presented in Section E.3.3.
  4. 4.
    Text-to-Text Generation: For Flan-T5-base and Flan-T5-large models fine-tuned with LoRA, the results are presented in Section E.3.4.
In Section E.3.1, E.3.2, E.3.3 and E.3.4, we compare the performance of different multi-task model fusion algorithms across various tasks. Pre-trained models’ performance, fine-tuned models’ performance, and traditional multi-task learning (MTL) methods are provided for reference. In Appendix D, we provide a detailed description of these fine-tuned single-task models.
We have the following key observations: (1) Multi-task model fusion usually outperforms pre-trained models, showing it can transfer knowledge from multiple single-task models to enhance performance. Pre-trained models lack task-specific knowledge as they are not fine-tuned for downstream tasks. (2) Adaptive method (AdaMerging) and MoE-based method perform best in multi-task model fusion, showing the effectiveness of adaptive merging and mixture-of-experts approaches. (3) The performance gap between multi-task model fusion and single-task fine-tuned models (STL) is larger for CLIP-ViT-B/32 compared to CLIP-ViT-L/14. This suggests that multi-task model fusion may be more beneficial for smaller models, as they have more room for improvement through knowledge transfer. (4) Traditional MTL outperforms most multi-task model fusion methods, which indicates that traditional MTL is still a strong baseline for multi-task learning, and there is room for improvement in multi-task model fusion algorithms.
E.3.1 Open-Vocabulary Image Classification Using CLIP-ViT Models
Figure 5: Radar charts comparing the performance of different deep model fusion methods across eight tasks using CLIP-ViT-B/32 and CLIP-ViT-L/14.
Table 8: Multi-task performance when merging CLIP-ViT-B/32 models on all eight tasks. Here underlined method names indicate methods that require test-time adaptation training, while italicized method names indicate model mixing approaches.
Method SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD AVG.
Reference Methods
Pre-trained 63.2 59.8 60.7 46.0 31.6 32.5 48.2 43.9 48.2
Individual Fine-tuned 75.0 78.3 95.2 99.0 97.3 98.9 99.6 79.7 90.3
Traditional MTL 72.3 76.6 92.2 97.9 95.5 97.7 99.3 77.7 88.6
Multi-Task Model Fusion Methods
Weight Averaging 65.4 62.6 70.8 76.9 64.5 54.9 86.3 50.9 66.5
Fisher Merging 66.7 64.0 72.2 91.6 69.0 64.3 83.5 53.7 70.6
RegMean 68.6 70.0 84.6 95.4 92.6 83.4 98.4 66.1 82.4
RegMean++ 69.3 70.5 86.7 96.1 94.1 90.4 99.0 68.7 84.4
Task Arithmetic 57.1 55.7 64.9 76.7 77.9 68.5 96.1 47.2 68.0
Ties-Merging 67.1 64.2 74.1 76.8 77.7 69.4 94.1 54.0 72.2
task-wise AdaMerging 58.6 56.9 69.8 82.4 70.3 58.9 97.2 55.3 68.7
layer-wise AdaMerging 67.9 71.3 83.5 92.7 87.4 92.9 98.2 67.0 82.6
WEMoE (Model Mixing) 73.7 76.8 93.4 98.2 96.8 98.2 99.6 76.6 89.2
SMILE (Model Mixing) 73.6 77.8 92.0 98.3 96.9 98.1 99.6 78.1 89.3
Table 9: Multi-task performance when merging CLIP-ViT-L/14 models on all eight tasks. Here underlined method names indicate methods that require test-time adaptation training, while italicized method names indicate model mixing approaches.
Method SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Avg.
Reference Methods
Pre-trained 68.3 77.8 71.0 58.9 58.4 50.6 76.4 55.5 64.6
Individual Fine-tuned 82.8 92.9 97.4 99.2 97.9 99.2 99.8 85.5 94.3
Traditional MTL 79.0 89.3 94.5 98.4 96.4 98.1 99.4 83.7 92.4
Multi-Task Model Fusion Methods
Weight Averaging 72.5 81.5 82.2 90.0 81.6 74.0 96.6 61.8 80.0
Fisher Merging 70.6 79.4 84.1 98.1 74.7 85.0 89.5 61.0 80.3
RegMean 76.9 89.8 93.0 97.5 96.3 94.1 98.7 77.0 90.4
RegMean++ 77.2 89.6 92.8 97.5 96.9 96.3 99.2 78.4 91.0
Task Arithmetic 72.0 79.0 80.5 86.0 87.5 83.5 98.0 58.8 80.7
Ties-Merging 74.7 83.3 86.4 91.3 89.7 85.2 97.8 63.9 84.0
task-wise AdaMerging 75.8 80.1 77.2 83.6 68.4 93.5 93.1 69.0 80.1
layer-wise AdaMerging 78.1 90.7 90.8 96.5 94.8 97.5 98.6 81.3 91.0
WEMoE (Model Mixing) 81.5 92.3 96.5 98.8 97.6 99.4 99.6 84.5 93.8
SMILE (Model Mixing) 81.9 92.3 95.5 99.1 98.0 98.9 99.7 83.6 93.6
We evaluate model fusion methods on open-vocabulary image classification using CLIP-ViT models fine-tuned on eight diverse image classification tasks spanning natural scenes, satellite imagery, handwritten digits, and texture recognition. Tables 8 and 9 present comprehensive results for CLIP-ViT-B/32 and CLIP-ViT-L/14 models respectively, while Figure 5 provides radar chart visualizations highlighting the performance patterns across tasks. The results demonstrate several key insights: (1) Model mixing methods (such as WEMoE and SMILE) consistently achieve the best performance, approaching traditional multi-task learning results with averages of 89.2-89.3% for ViT-B/32 and 93.6-93.8% for ViT-L/14. (2) Layer-wise adaptive merging (AdaMerging) significantly outperforms task-wise variants, highlighting the importance of fine-grained parameter-level fusion strategies. (3) Simple parameter averaging and Fisher merging provide strong baselines but fall short of more sophisticated approaches like RegMean, which leverages data alignment techniques. (4) The performance gap between fusion methods and individual fine-tuned models is more pronounced for the smaller ViT-B/32 (90.3% vs 89.3%) compared to ViT-L/14 (94.3% vs 93.8%), suggesting that larger models may be more amenable to effective knowledge combination through fusion techniques. This phenomenon has also been discussed in the literature on large language model merging (yadav2024matters).
E.3.2 Scene Understanding Using ResNet-50 Models
@@
Method SEGMENTATION DEPTH ESTIMATION NORMAL
Single-Task Learning
Segmentation 52.0 73.8 242.8 88.7 82.8
Depth Estimation 2.3 6.2 42.5 17.7 82.8
Normal 2.0 4.9 264.0 98.1 24.7
Multi-Task Model Fusion Methods
Weight Averaging 39.0 67.0 55.1 22.7 30.4
Task Arithmetic ($\lambda=0.3$) 33.6 63.3 56.3 23.2 31.3
Ties-Merging ($\lambda=0.3$) 36.3 61.7 60.5 24.5 33.1
Figure 6 illustrates the multi-task model fusion problem in scene understanding, where three specialized ResNet-50 models trained individually on segmentation, depth estimation, and normal estimation tasks are merged at the backbone level while preserving task-specific heads. Table 10 presents the results. The results show that single-task ResNet50 models excel only at their own objectives but fail badly on other tasks, highlighting strong specialization without transferability. In contrast, model fusion produces balanced performance across tasks: for example, while fusion methods sacrifice some segmentation accuracy compared to the segmentation expert, they dramatically improve depth and normal performance relative to non-specialists.
Table 11: Multi-task performance of GPT-2 models on seven text classification tasks.
@@
Multi-Task Deep Model Fusion Method
Simple Average 55.0 55.1 51.0 57.6 76.7 44.8 52.5 56.1
Fisher Merging 54.8 58.0 39.5 63.3 81.5 49.1 64.7 58.7
RegMean 61.7 70.4 65.4 69.7 78.8 56.0 79.7 68.8
Task Arithmetic 68.7 68.6 69.6 70.5 81.8 47.3 83.6 70.0
Ties-Merging 68.4 71.4 68.4 69.6 82.4 47.7 81.8 70.0
E.3.3 Text Classification Using GPT-2 Models
We evaluate multi-task model fusion methods on natural language understanding tasks using GPT-2 models fine-tuned on seven diverse text classification tasks from the GLUE benchmark (wang2018glue). Table 11 presents the results.
E.3.4 Text-to-Text Generation Using Flan-T5 Models
Table 12: Experimental results of merging Flan-T5-base (LoRA fine-tuned) models on all eight tasks. Here underlined method names indicate methods that require test-time adaptation training, while italicized method names indicate model mixing approaches.
@@ E.3.4 Text-to-Text Generation Using Flan-T5 Models
Weight Averaging 69.7 59.7 78.9 90.1 83.8 80.5 91.2 72.0 78.2
Task Arithmetic 68.8 55.2 78.7 89.8 83.7 79.1 91.5 72.4 77.4
Ties-Merging 68.3 56.3 79.4 89.8 83.7 79.4 91.6 71.2 77.5
Layer-wise AdaMerging 69.1 60.3 78.4 90.0 83.6 79.1 91.6 74.1 78.3
SMILE (Model Mixing) 69.3 82.9 83.8 90.6 83.9 83.4 93.1 85.1 84.0
@@
Weight Averaging 74.6 84.3 84.1 92.8 86.3 87.4 94.8 88.0 86.5
Task Arithmetic 76.9 85.4 85.3 93.9 85.8 88.1 95.2 87.8 87.3
Ties-Merging 77.1 85.1 86.3 93.9 86.0 87.7 95.1 88.0 87.4
Layer-wise AdaMerging 76.7 87.6 84.8 93.8 85.9 88.1 95.2 88.6 87.6
We extend our evaluation to text-to-text generation tasks using Flan-T5 models fine-tuned with LoRA on eight tasks from the GLUE benchmark (wang2018glue), including both classification and regression tasks. Tables 12 and 13 present results for Flan-T5-base and Flan-T5-large models respectively, demonstrating how model fusion performs in the parameter-efficient fine-tuning paradigm. The LoRA adapters are merged into the base models before fusion, creating full fine-tuned models that are then subject to various fusion techniques. Notably, while Flan-T5-base fusion methods show modest improvements over simple averaging (78.2% vs. individual fine-tuned performance of 84.6%), the SMILE model mixing method achieves remarkable performance (84.0%), nearly matching the individual fine-tuned baseline. For Flan-T5-large, all fusion methods demonstrate stronger performance, with the best methods (Task Arithmetic, Ties-Merging, Layer-wise AdaMerging) achieving 87.3-87.6% compared to 89.6% for individual models, suggesting that larger models provide more robust representations for effective knowledge combination across tasks. This trend aligns with observations from the vision experiments, reinforcing the hypothesis that model scale facilitates more effective parameter-space fusion.
E.4 Generalization and Robustness Evaluation
To further assess the generalization and robustness of multi-task model fusion algorithms, we conduct experiments on unseen tasks and corrupted test sets (or out-of-distribution test sets).
From the experimental results in Section E.4.1 and E.4.2, we have the following key observations: (1) The performance of all multi-task model fusion methods on unseen tasks is generally lower than their performance on seen tasks. This is expected, as the models being fused are not explicitly trained on the unseen tasks. (2) A negative transfer is observed in Table 15 on the RESISC45 dataset, where the merged models exhibit lower accuracy compared to the pre-trained model. The performance of all multi-task model fusion methods on RESISC45 is lower than the pre-trained model, indicating that the knowledge transferred from the seen tasks may not be beneficial or even harmful to this specific unseen task. (3) The performance of all methods drops significantly on certain types of corruptions, such as pixelation and impulse noise. This highlights the challenge of maintaining robustness under severe distribution shifts and the need for further research in this direction. (4) When the test distribution is corrupted, adaptive methods may overfit to certain tasks, leading to a decrease in overall performance. This suggests that adaptive methods may need to be further regularized to improve generalization and robustness.
E.4.1 Generalization Experiments
@@
Table 14: Generalization results when merging ViT-B/32 models on six tasks (SUN397, Cars, RESISC45, DTD, SVHN, GTSRB) and evaluating on two unseen tasks (MNIST, EuroSAT).
Method Seen Tasks (ACC) Unseen Tasks (ACC)
SUN397 Cars RESISC45 DTD SVHN GTSRB Avg. MNIST EuroSAT Avg.
Pre-trained 63.2 59.9 60.6 43.9 23.5 30.4 46.9 47.6 45.6 46.6
Fisher Merging 65.5 67.2 78.2 57.6 84.2 75.9 71.4 71.8 49.4 60.6
RegMean 68.7 70.0 86.5 65.9 93.9 86.7 78.6 82.2 49.3 65.7
Task Arithmetic 64.3 63.0 73.2 54.9 84.7 79.5 69.9 75.5 42.6 59.1
Ties-Merging 68.3 65.5 76.9 54.9 75.4 72.0 68.9 73.1 47.3 60.2
AdaMerging 68.4 71.9 87.9 69.1 92.2 93.8 80.5 77.7 47.3 62.5
WEMoE 75.4 77.5 94.3 77.0 96.8 98.7 86.6 78.3 44.0 61.1
Table 15: Generalization results when merging ViT-B/32 models on six tasks (SUN397, Cars, GTSRB, EuroSAT, DTD, MNIST) and evaluating on two unseen tasks (RESISC45, SVHN).
Method Seen Tasks (ACC) Unseen Tasks (ACC)
@@
Fisher Merging 68.1 67.4 67.2 86.4 58.6 81.6 71.5 60.2 42.5 51.3
RegMean 69.4 70.5 86.9 97.0 67.1 98.3 81.5 50.2 51.5 50.8
Task Arithmetic 65.2 63.6 76.1 87.1 56.4 94.2 73.8 52.4 45.2 48.8
Ties-Merging 68.2 65.9 70.0 81.2 56.0 89.0 71.7 60.3 47.3 53.8
AdaMerging 69.8 72.4 95.5 95.1 70.7 98.1 83.6 48.7 60.7 54.7
WEMoE 74.3 78.1 98.8 98.7 75.1 99.5 87.4 47.3 51.3 49.3
For the generalization experiments, we assess the performance of multi-task deep model fusion algorithms on two unseen tasks after merging ViT-B/32 models trained on six tasks. The performance of various multi-task model fusion methods, including Fisher Merging (matena2022merging), RegMean (jin2022dataless), Task Arithmetic (ilharco2022editing), Ties-Merging (yadav2023resolving), AdaMerging (yang2023adamerging), and WEMoE (tang2024merging), is compared across both the seen tasks and unseen tasks. Specifically, we conduct two sets of generalization experiments using the CLIP-ViT-B/32 models: